# Large Language Models in Argument Mining: A Survey

### v1

@@ Abstract
This work is in progress. The content and references are subject to change.
@@ 1 Introduction
Argument Mining (AM) is a fast-growing subfield of Natural Language Processing (NLP) concerned with the automatic identification and extraction of argumentative structures—claims, premises, and the relations between them—from textual discourse (Lawrence & Reed, 2019; Patel, 2024). Such structure is central to understanding not just what opinions are expressed but why, enabling applications ranging from debate moderation and fact checking to legal analytics and scientific discourse analysis. Indeed, by transforming unstructured text into structured argument representations, AM supports reasoning-centered applications that go far beyond traditional sentiment analysis, which can determine opinions but not their underlying rationale (Lawrence & Reed, 2019).
@@
Despite its clear value, manual argument annotation remains labor-intensive and inherently unscalable: crafting detailed argument maps can consume hundreds or even thousands of hours for relatively short texts. To cope with the volume of data generated online, research in Argument Mining (AM) has accelerated significantly since around 2013, marked by the launch of dedicated workshops at ACL and themed tutorials at top-tier conferences including IJCAI, ACL, and ESSLLI (Cabrio & Villata, 2018). Foundational surveys—most notably those by Lawrence & Reed (2019), which comprehensively cataloged AM techniques, datasets, and challenges (Lawrence & Reed, 2019); Peldszus & Stede (2013), who systematically bridged classical argument-diagramming methods with machine-readable annotation pipelines (Peldszus & Stede, 2013); and Lytos et al. (2019), who documented AM’s adaptation to less formal, user-generated text (Lytos et al., 2019)—have provided essential context for early developments.
@@
Building on these foundations, subsequent studies have enriched the field with domain- and method-specific insights. Cabrio & Villata (2018) offered a data-driven reflection on the field’s early corpus and model evolution , while targeted surveys explored annotation practices, tool ecosystems, and argument quality assessment. For example, Patel (2024) traces the integration of traditional machine learning with emerging neural techniques; Schaefer & Stede (2021) focus on Twitter; Hua et al. (2019) and Fromm et al. (2021) examine peer-review text; and Guida et al. (2025) assess the performance of LLMs on enforcing topic-specific argument tasks in online comments (Altemeyer et al., 2025; Chen et al., 2024).
@@
However, despite this breadth, most existing reviews remain anchored in the pre‑LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, and reconnects these technical developments to foundational argumentation theory and annotation practice.
@@
Building on this foundation, the rest of the survey unfolds as follows.Section 2 catalogs datasets and benchmarks—emphasizing recent large, cross-domain, and multimodal resources suited for LLM evaluation. Section 3 develops a taxonomy of key AM subtasks—claim detection, premise identification, relation parsing, quality assessment, and summarization—exploring how LLM techniques like prompting, chain‑of‑thought, few‑shot learning, and retrieval‑augmented generation reshape each. Section 4 introduces model architectures and key LLM-era techniques, detailing encoder–decoder designs, prompting strategies, fine-tuning approaches, and multimodal integration relevant to argument mining. Section 5 outlines evaluation methodologies, including metrics, human-in-the-loop assessments, and benchmarking practices tailored to AM. Section 6 delves into ongoing challenges—such as long-context reasoning, domain transfer, interpretability, annotation bottlenecks, and computational efficiency—while highlighting emerging research trends. Finally, Section 7 presents our conclusions and proposes a forward-looking roadmap for advancing LLM-powered computational argumentation.
@@ 2.1 Capturing Argumentation in Dialogue
Moving towards a more granular analysis of dialogue mechanics, the CAMAL (Conversational Argument Move Analysis) dataset introduces a multi-label framework for annotating ”conversational moves” Lai et al. (2024). This approach moves beyond the traditional support/attack dichotomy to capture the multifaceted functions of utterances, such as questioning, clarifying, or challenging a premise. Such a fine-grained annotation scheme more accurately reflects the fluidity of natural dialogue and supports the development of models for more sophisticated dialogical systems. The challenge of parsing multi-party dialogue is addressed by the work of Mao et al., who developed a new argumentative dataset alongside a hierarchical graph fusion network Mao et al. (2024). Their contribution underscores the necessity of understanding global discourse structure for accurately resolving local argumentative relations in complex, multi-participant settings. Reinforcing this theme, the work of Mezza et al. demonstrates that exploiting dialogue acts and contextual information significantly improves the identification of argumentative relations in online debates Mezza et al. (2024).
@@ 2.3 Argument Structure
Concurrent with the expansion of scope, there is a clear trend towards deepening the analysis of core argumentative concepts and applying them in highly specialized domains. The legal field, with its inherently structured and argumentative nature, remains a key area of application. The dataset of annotated European Court of Human Rights (ECHR) cases presented by Mumford et al. offers detailed insights into the intricacies of legal reasoning Mumford et al. (2024). Similarly, in the medical domain, the CasiMedicos-Arg dataset annotates explanatory argumentative structures within medical question-answering contexts, linking argument mining directly to the critical need for explainability in clinical AI Sviridova et al. (2024a; b).
@@ 2.4 Data Creation and Evaluation in LLM Era
The advent of Large Language Models (LLMs) has introduced a new dynamic into the creation and use of argumentative datasets. On one hand, LLMs are being leveraged as tools to overcome data scarcity. The work of de Souza and Dragoni, for instance, explores the use of LLMs to augment persuasive argument datasets, offering a scalable method for generating new training instances de Souza & Dragoni (2024). On the other hand, the argumentative capabilities of LLMs are themselves becoming a central object of study. Researchers are now creating datasets specifically designed to probe and evaluate the quality of arguments generated by models like ChatGPT. The work by Rocha et al. provides a methodology and dataset to assess ”good, bad, and ugly” machine-generated arguments Rocha et al. (2024), while Djouvas et al. present a comparative study and dataset for argument strength classification that critically evaluates the performance of LLMs against other models Djouvas et al. (2024). This reflexive turn, where LLMs are used both to create and to be evaluated by data, marks a significant new phase in the field, enabling a more nuanced understanding of the strengths and limitations of current generative models in the context of complex human reasoning.
@@ 3.1.2 Domain-Specialized Detection Applications
Healthcare & Biomedical: This critical domain combines fraud detection Gao et al. (2019), privacy-preserving blockchain verification Mohan & Praveen (2019), and evolutionary optimization Tubishat et al. (2025). Biomedical applications focus on Twitter-based health claims Wührl & Klinger (2021) and contradictory research detection Yazi et al. (2021). Recent systematic reviews du Preez et al. (2025) evaluate machine learning efficacy across healthcare fraud patterns.
@@ 3.1.3 Contextual Application Scenarios
Political Fact-Checking: This high-stakes domain employs innovative fairness techniques like actor masking Dayanik & Padó (2020) and positive unlabeled learning Wright & Augenstein (2020). Advanced verification frameworks Jafari & Allan (2024) improve robustness against adversarial political messaging, with rumor detection systems Lin et al. (2021) using claim-guided hierarchical attention networks to combat misinformation.
@@ 3.1.4 Cross-Lingual Methodologies
Multilingual detection addresses global misinformation through adapter fusion techniques Schlicht et al. (2023) and LLM-based approaches Vykopal et al. (2025). Cross-lingual transfer learning Panchendrarajan & Zubiaga (2025) enables knowledge sharing between languages, while transformer-based multilingual fact claiming Böck et al. (2021) demonstrates effectiveness across diverse linguistic landscapes.
@@ 3.1.5 Verification and Anomaly Detection
Claim Verification: Sophisticated validation frameworks Irnawan et al. (2025) assess claim authenticity through evidence retrieval and reasoning chains. Robust verification techniques Jafari & Allan (2024) address adversarial examples and out-of-distribution claims in high-risk domains.
@@ 3.2.1 LLMs in Evidence Detection
Several studies have focused on employing pretrained language models for automated evidence extraction. Dementieva and Panchenko Dementieva & Panchenko (2021; 2020) demonstrated the effectiveness of cross-lingual evidence in bolstering monolingual fake news detection. Elaraby and Litman Elaraby & Litman (2021) employed self-trained pretrained language models specifically optimized for evidence detection, achieving notable improvements. Graph-based techniques also constitute a vital approach within evidence detection. Xu et al. Xu et al. (2022a; b) introduced an evidence-aware fake news detection method utilizing graph neural networks to mine fine-grained semantics. Moreover, Guo et al. Guo et al. (2023) developed an interpretable fake news detection model leveraging graph-based evidence, enhancing interpretability and accuracy. Interactive and dynamic evidence retrieval methodologies have also emerged. Liao et al. Liao et al. (2023a; b) proposed a multi-step evidence retrieval enhancement framework (MUSER), significantly improving accuracy by iterative evidence refinement. Furthermore, Wu et al. Wu et al. (2024) developed a unified evidence enhancement inference framework that dynamically adapts to incoming information.
@@ 3.2.2 Multilingual Evidence Detection
A significant trend involves evidence extraction across multiple sources or languages. Dementieva et al. Dementieva et al. (2023) proposed a multilingual approach named Multiverse, integrating evidence from diverse linguistic contexts. Hammouchi and Ghogho Hammouchi & Ghogho (2022) similarly developed an evidence-aware multilingual fake news detection system, underscoring the importance of linguistic diversity in robust misinformation detection.
@@ 3.2.4 Multil-modal Evidence Detection
Multi-modal rumor detection has seen significant enhancements through dual evidence mechanisms. Huang et al. Huang et al. (2025) proposed dual evidence enhancement alongside text-image similarity awareness, significantly improving misinformation detection efficacy. Similarly, Wu and Cao Wu & Cao (2024) developed an evidence-aware multimodal framework tailored specifically for Chinese social media, effectively capturing the nuances inherent in multimodal data.
@@ 3.3 Stance Detection
Recent efforts have produced a variety of stance detection datasets to support research in this area. For instance, Niu et al. (2025a; b) introduce C-MTCSD, a Chinese multi-turn conversational stance detection dataset, which captures nuanced stance evolution. Gyawali et al. (2024) contribute GunStance, a dataset focusing on gun regulation debates. These datasets facilitate stance classification in diverse linguistic, political, and conversational contexts.
@@
Multilingual stance detection remains a challenge. Charfi et al. (2024) present a multi-dialect Arabic stance corpus and cross-domain benchmarks. Works like Alghaslan & Almutairy (2024); Shukla et al. (2024) leverage LLM fine-tuning and ensembling for Arabic stance detection tasks.
@@
The promise of zero-shot stance detection has spurred further innovation. Ma et al. (2024a; b) propose Chain-of-Stance, where LLMs perform chain-of-thought style reasoning. Guo et al. (2024) incorporate structured external knowledge from LLMs to improve generalization. Zhang et al. (2024) integrate commonsense knowledge through adversarial learning.
@@
Acknowledging LLM bias, Li et al. (2025; 2024a) propose counterfactual augmented calibration strategies to reduce stance-related biases in LLM outputs.
@@
LLM-generated synthetic datasets are also explored. For example, Wagner et al. (2025) show that LLM-synthesized stance data can outperform traditional fine-tuning when training low-resource classifiers. Similarly, Fraile-Hernandez & Peñas (2024) demonstrate the use of generative LLMs in political activism scenarios.
@@
Finally, de Landa & Agerri (2025) present language-agnostic methods using interaction-based embeddings and multilingual LLMs to promote cross-lingual stance transfer.
@@ 3.4 Argument Summarization
Early work such as ConvoSumm(Fabbri et al., 2021) laid the foundation for integrating argument mining with abstractive summarization in dialogues, demonstrating the benefits of identifying argumentative components in conversation summarization. Building on this, several works have extended the scope of argument summarization to new domains such as law(Elaraby et al., 2023; Elaraby & Litman, 2022; Xu & Ashley, 2023; Steffes & Rataj, 2022), where legal documents are often long and require nuanced treatment of argumentative structure. The ArgLegalSumm model(Elaraby & Litman, 2022) proposed incorporating argument mining to enhance legal document summaries, and was further extended with reranking modules to improve informativeness and coherence(Elaraby et al., 2023).
In terms of modeling techniques, Shan and Lu(Shan & Lu, 2025) applied multi-task learning with argument relation-aware contrastive learning to better capture opinion divergence in debates. Similarly, Syed et al.(Syed et al., 2023) introduced frame-oriented summarization to capture pragmatic aspects of arguments in discussions. Other efforts, such as by Meer et al.(van der Meer et al., 2024), analyzed the diversity of generated summaries, revealing trade-offs between informativeness and representation of varied perspectives.

### v2

@@ 1 Introduction
Argument Mining (AM) is a fast-growing subfield of Natural Language Processing (NLP) concerned with the automatic identification and extraction of argumentative structures—claims, premises, and the relations between them—from textual discourse (Lawrence and Reed, 2019; Patel, 2024). Such structure is central to understanding not just what opinions are expressed but why, enabling applications ranging from debate moderation and fact checking to legal analytics and scientific discourse analysis. Indeed, by transforming unstructured text into structured argument representations, AM supports reasoning-centered applications that go far beyond traditional sentiment analysis, which can determine opinions but not their underlying rationale (Lawrence and Reed, 2019).
@@
Despite its clear value, manual argument annotation remains labor-intensive and inherently unscalable: crafting detailed argument maps can consume hundreds or even thousands of hours for relatively short texts. To cope with the volume of data generated online, research in Argument Mining (AM) has accelerated significantly since around 2013, marked by the launch of dedicated workshops at ACL and themed tutorials at top-tier conferences including IJCAI, ACL, and ESSLLI (Cabrio and Villata, 2018). Foundational surveys—most notably those by Lawrence and Reed (2019), which comprehensively cataloged AM techniques, datasets, and challenges (Lawrence and Reed, 2019); Peldszus and Stede (2013), who systematically bridged classical argument-diagramming methods with machine-readable annotation pipelines (Peldszus and Stede, 2013); and Lytos et al. (2019), who documented AM’s adaptation to less formal, user-generated text (Lytos et al., 2019)—have provided essential context for early developments.
@@
Building on these foundations, subsequent studies have enriched the field with domain- and method-specific insights. Cabrio and Villata (2018) offered a data-driven reflection on the field’s early corpus and model evolution , while targeted surveys explored annotation practices, tool ecosystems, and argument quality assessment. For example, Patel (2024) traces the integration of traditional machine learning with emerging neural techniques; Schaefer and Stede (2021) focus on Twitter; Hua et al. (2019) and Fromm et al. (2021) examine peer-review text; and Guida et al. (2025) assess the performance of LLMs on enforcing topic-specific argument tasks in online comments (Altemeyer et al., 2025; Chen et al., 2024).
@@
However, despite this breadth, most existing reviews remain anchored in the pre-LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, and reconnects these technical developments to foundational argumentation theory and annotation practice.
@@
Building on this foundation, the rest of the survey unfolds as follows.Section 2 catalogs datasets and benchmarks—emphasizing recent large, cross-domain, and multimodal resources suited for LLM evaluation. Section 3 develops a taxonomy of key AM subtasks—claim detection, premise identification, relation parsing, quality assessment, and summarization—exploring how LLM techniques like prompting, chain-of-thought, few-shot learning, and retrieval-augmented generation reshape each. Section 4 introduces model architectures and key LLM-era techniques, detailing encoder–decoder designs, prompting strategies, fine-tuning approaches, and multimodal integration relevant to argument mining. Section 5 outlines evaluation methodologies, including metrics, human-in-the-loop assessments, and benchmarking practices tailored to AM. Section 6 delves into ongoing challenges—such as long-context reasoning, domain transfer, interpretability, annotation bottlenecks, and computational efficiency—while highlighting emerging research trends. Finally, Section 7 presents our conclusions and proposes a forward-looking roadmap for advancing LLM-powered computational argumentation.
@@ 2.1 Capturing Argumentation in Dialogue
Moving towards a more granular analysis of dialogue mechanics, the CAMAL (Conversational Argument Move Analysis) dataset introduces a multi-label framework for annotating "conversational moves" Lai et al. (2024). This approach moves beyond the traditional support/attack dichotomy to capture the multifaceted functions of utterances, such as questioning, clarifying, or challenging a premise. Such a fine-grained annotation scheme more accurately reflects the fluidity of natural dialogue and supports the development of models for more sophisticated dialogical systems. The challenge of parsing multi-party dialogue is addressed by the work of Mao et al., who developed a new argumentative dataset alongside a hierarchical graph fusion network Mao et al. (2024). Their contribution underscores the necessity of understanding global discourse structure for accurately resolving local argumentative relations in complex, multi-participant settings. Reinforcing this theme, the work of Mezza et al. demonstrates that exploiting dialogue acts and contextual information significantly improves the identification of argumentative relations in online debates Mezza et al. (2024).
@@ 2.3 Argument Structure
Concurrent with the expansion of scope, there is a clear trend towards deepening the analysis of core argumentative concepts and applying them in highly specialized domains. The legal field, with its inherently structured and argumentative nature, remains a key area of application. The dataset of annotated European Court of Human Rights (ECHR) cases presented by Mumford et al. offers detailed insights into the intricacies of legal reasoning Mumford et al. (2024). Similarly, in the medical domain, the CasiMedicos-Arg dataset annotates explanatory argumentative structures within medical question-answering contexts, linking argument mining directly to the critical need for explainability in clinical AI Sviridova et al. (2024a, b).
@@ 2.4 Data Creation and Evaluation in LLM Era
The advent of Large Language Models (LLMs) has introduced a new dynamic into the creation and use of argumentative datasets. On one hand, LLMs are being leveraged as tools to overcome data scarcity. The work of de Souza and Dragoni, for instance, explores the use of LLMs to augment persuasive argument datasets, offering a scalable method for generating new training instances de Souza and Dragoni (2024). On the other hand, the argumentative capabilities of LLMs are themselves becoming a central object of study. Researchers are now creating datasets specifically designed to probe and evaluate the quality of arguments generated by models like ChatGPT. The work by Rocha et al. provides a methodology and dataset to assess "good, bad, and ugly" machine-generated arguments Rocha et al. (2024), while Djouvas et al. present a comparative study and dataset for argument strength classification that critically evaluates the performance of LLMs against other models Djouvas et al. (2024). This reflexive turn, where LLMs are used both to create and to be evaluated by data, marks a significant new phase in the field, enabling a more nuanced understanding of the strengths and limitations of current generative models in the context of complex human reasoning.
@@ 3.1.2 Domain-Specialized Detection Applications
Healthcare & Biomedical: This critical domain combines fraud detection Gao et al. (2019), privacy-preserving blockchain verification Mohan and Praveen (2019), and evolutionary optimization Tubishat et al. (2025). Biomedical applications focus on Twitter-based health claims Wührl and Klinger (2021) and contradictory research detection Yazi et al. (2021). Recent systematic reviews du Preez et al. (2025) evaluate machine learning efficacy across healthcare fraud patterns.
@@ 3.1.3 Contextual Application Scenarios
Political Fact-Checking: This high-stakes domain employs innovative fairness techniques like actor masking Dayanik and Padó (2020) and positive unlabeled learning Wright and Augenstein (2020). Advanced verification frameworks Jafari and Allan (2024) improve robustness against adversarial political messaging, with rumor detection systems Lin et al. (2021) using claim-guided hierarchical attention networks to combat misinformation.
@@ 3.1.4 Cross-Lingual Methodologies
Multilingual detection addresses global misinformation through adapter fusion techniques Schlicht et al. (2023) and LLM-based approaches Vykopal et al. (2025). Cross-lingual transfer learning Panchendrarajan and Zubiaga (2025) enables knowledge sharing between languages, while transformer-based multilingual fact claiming Böck et al. (2021) demonstrates effectiveness across diverse linguistic landscapes.
@@ 3.1.5 Verification and Anomaly Detection
Claim Verification: Sophisticated validation frameworks Irnawan et al. (2025) assess claim authenticity through evidence retrieval and reasoning chains. Robust verification techniques Jafari and Allan (2024) address adversarial examples and out-of-distribution claims in high-risk domains.
@@ 3.2.1 LLMs in Evidence Detection
Several studies have focused on employing pretrained language models for automated evidence extraction. Dementieva and Panchenko Dementieva and Panchenko (2021, 2020) demonstrated the effectiveness of cross-lingual evidence in bolstering monolingual fake news detection. Elaraby and Litman Elaraby and Litman (2021) employed self-trained pretrained language models specifically optimized for evidence detection, achieving notable improvements. Graph-based techniques also constitute a vital approach within evidence detection. Xu et al. Xu et al. (2022a, b) introduced an evidence-aware fake news detection method utilizing graph neural networks to mine fine-grained semantics. Moreover, Guo et al. Guo et al. (2023) developed an interpretable fake news detection model leveraging graph-based evidence, enhancing interpretability and accuracy. Interactive and dynamic evidence retrieval methodologies have also emerged. Liao et al. Liao et al. (2023a, b) proposed a multi-step evidence retrieval enhancement framework (MUSER), significantly improving accuracy by iterative evidence refinement. Furthermore, Wu et al. Wu et al. (2024) developed a unified evidence enhancement inference framework that dynamically adapts to incoming information.
@@ 3.2.2 Multilingual Evidence Detection
A significant trend involves evidence extraction across multiple sources or languages. Dementieva et al. Dementieva et al. (2023) proposed a multilingual approach named Multiverse, integrating evidence from diverse linguistic contexts. Hammouchi and Ghogho Hammouchi and Ghogho (2022) similarly developed an evidence-aware multilingual fake news detection system, underscoring the importance of linguistic diversity in robust misinformation detection.
@@ 3.2.4 Multil-modal Evidence Detection
Multi-modal rumor detection has seen significant enhancements through dual evidence mechanisms. Huang et al. Huang et al. (2025) proposed dual evidence enhancement alongside text-image similarity awareness, significantly improving misinformation detection efficacy. Similarly, Wu and Cao Wu and Cao (2024) developed an evidence-aware multimodal framework tailored specifically for Chinese social media, effectively capturing the nuances inherent in multimodal data.
@@ 3.3 Stance Detection
Recent efforts have produced a variety of stance detection datasets to support research in this area. For instance, Niu et al. (2025a, b) introduce C-MTCSD, a Chinese multi-turn conversational stance detection dataset, which captures nuanced stance evolution. Gyawali et al. (2024) contribute GunStance, a dataset focusing on gun regulation debates. These datasets facilitate stance classification in diverse linguistic, political, and conversational contexts.
@@
Multilingual stance detection remains a challenge. Charfi et al. (2024) present a multi-dialect Arabic stance corpus and cross-domain benchmarks. Works like Alghaslan and Almutairy (2024); Shukla et al. (2024) leverage LLM fine-tuning and ensembling for Arabic stance detection tasks.
@@
The promise of zero-shot stance detection has spurred further innovation. Ma et al. (2024a, b) propose Chain-of-Stance, where LLMs perform chain-of-thought style reasoning. Guo et al. (2024) incorporate structured external knowledge from LLMs to improve generalization. Zhang et al. (2024) integrate commonsense knowledge through adversarial learning.
@@
Acknowledging LLM bias, Li et al. (2025, 2024a) propose counterfactual augmented calibration strategies to reduce stance-related biases in LLM outputs.
@@
LLM-generated synthetic datasets are also explored. For example, Wagner et al. (2025) show that LLM-synthesized stance data can outperform traditional fine-tuning when training low-resource classifiers. Similarly, Fraile-Hernandez and Peñas (2024) demonstrate the use of generative LLMs in political activism scenarios.
@@
Finally, de Landa and Agerri (2025) present language-agnostic methods using interaction-based embeddings and multilingual LLMs to promote cross-lingual stance transfer.
@@ 3.4 Argument Summarization
Early work such as ConvoSumm(Fabbri et al., 2021) laid the foundation for integrating argument mining with abstractive summarization in dialogues, demonstrating the benefits of identifying argumentative components in conversation summarization. Building on this, several works have extended the scope of argument summarization to new domains such as law(Elaraby et al., 2023; Elaraby and Litman, 2022; Xu and Ashley, 2023; Steffes and Rataj, 2022), where legal documents are often long and require nuanced treatment of argumentative structure. The ArgLegalSumm model(Elaraby and Litman, 2022) proposed incorporating argument mining to enhance legal document summaries, and was further extended with reranking modules to improve informativeness and coherence(Elaraby et al., 2023).
In terms of modeling techniques, Shan and Lu(Shan and Lu, 2025) applied multi-task learning with argument relation-aware contrastive learning to better capture opinion divergence in debates. Similarly, Syed et al.(Syed et al., 2023) introduced frame-oriented summarization to capture pragmatic aspects of arguments in discussions. Other efforts, such as by Meer et al.(van der Meer et al., 2024), analyzed the diversity of generated summaries, revealing trade-offs between informativeness and representation of varied perspectives.